# scores: A Python package for verifying and evaluating models and predictions with xarray and pandas

### v1

@@
scores: A Python package for verifying and evaluating models and predictions with xarray and pandas
Tennessee Leeuwenburg
@@ 1 Summary
scores is a Python package containing mathematical functions for the verification, evaluation and optimisation of forecasts, predictions or models. It primarily supports the geoscience communities; in particular, the meteorological, climatological and oceanographic communities. In addition to supporting the Earth system science communities, it also has wide potential application in machine learning and other domains such as economics.
scores not only includes common scores (e.g. Mean Absolute Error), it also includes novel scores not commonly found elsewhere (e.g. FIxed Risk Multicategorical (FIRM) score, Flip-Flop Index), complex scores (e.g. threshold-weighted continuous ranked probability score), and statistical tests (such as the Diebold Mariano test). It also contains isotonic regression which is becoming an increasingly important tool in forecast verification and can be used to generate stable reliability diagrams. Additionally, it provides pre-processing tools for preparing data for scores in a variety of formats including cumulative distribution functions (CDF). At the time of writing, scores includes over 50 metrics, statistical techniques and data processing tools.
@@
scores primarily supports xarray datatypes for Earth system data, allowing it to work with NetCDF4, HDF5, Zarr and GRIB data sources among others. scores uses Dask for scaling and performance. It has expanding support for pandas.
The software repository can be found at https://github.com/nci/scores/.
@@ 2 Statement of Need
The purpose of this software is (a) to mathematically verify and validate models and predictions and (b) to foster research into new scores and metrics.
@@ 2.1 Key Benefits of scores
In order to meet the needs of researchers and other users, scores provides the following key benefits.
Data Handling
  • Works with n-dimensional data (e.g., geospatial, vertical and temporal dimensions) for both point-based and gridded data. scores can effectively handle the dimensionality, data size and data structures commonly used for:
    • gridded Earth system data (e.g. numerical weather prediction models)
    • tabular, point, latitude/longitude or site-based data (e.g. forecasts for specific locations).
@@
  • Supports xarray (Hoyer & Hamman, 2017) datatypes, and works with NetCDF4 (Unidata, 2024), HDF5 (The HDF Group & Koziol, 2020), Zarr (Miles et al., 2020) and GRIB (World Meteorological Organization, 2024) data sources among others.
Usability
@@
  • Novel scores not commonly found elsewhere (e.g. FIRM (Taggart, Loveday, & Griffiths, 2022), Flip-Flop Index (Griffiths, Foley, Ioannou, & Leeuwenburg, 2019; Griffiths, Loveday, Price, Foley, & McKelvie, 2021)).
@@
  • Expanding support for pandas (McKinney, 2010; The pandas development team, 2024).
@@ 2.2 Metrics, Statistical Techniques and Data Processing Tools Included in scores
Continuous Scores for evaluating single-valued continuous forecasts. Mean Absolute Error (MAE), Mean Squared Error (MSE), Root Mean Squared Error (RMSE), Additive Bias, Multiplicative Bias, Pearson’s Correlation Coefficient, Flip-Flop Index (Griffiths et al., 2019, 2021), Quantile Loss, Murphy Score (Ehm, Gneiting, Jordan, & Krüger, 2016).
Probability Scores for evaluating forecasts that are expressed as predictive distributions, ensembles, and probabilities of binary events. Brier Score (Brier, 1950), Continuous Ranked Probability Score (CRPS) for Cumulative Distribution Functions (CDFs) (including threshold-weighting, see Gneiting & Ranjan (2011)), CRPS for ensembles (Ferro, 2013; Gneiting & Raftery, 2007), Receiver Operating Characteristic (ROC), Isotonic Regression (reliability diagrams) (Dimitriadis, Gneiting, & Jordan, 2021).
Categorical Scores for evaluating forecasts of categories. Probability of Detection (POD), Probability of False Detection (POFD), False Alarm Ratio (FAR), Success Ratio, Accuracy, Peirce’s Skill Score (Peirce, 1884), Critical Success Index (CSI), Gilbert Skill Score (Gilbert, 1884), Heidke Skill Score, Odds Ratio, Odds Ratio Skill Score, F1 Score, Symmetric Extremal Dependence Index (Ferro & Stephenson, 2011), FIxed Risk Multicategorical (FIRM) Score (Taggart et al., 2022).
@@
Statistical Tests Tools to conduct statistical tests and generate confidence intervals. Diebold-Mariano (Diebold & Mariano, 1995) with both the Harvey, Leybourne, & Newbold (1997) and Hering & Genton (2011) modifications.
@@ 2.3 Use in Academic Work
In 2015, the Australian Bureau of Meteorology began developing a new verification system called Jive, which became operational in 2022. For a description of Jive see Loveday, Griffiths, et al. (2024). The Jive verification metrics have been used to support several publications (Foley & Loveday, 2020; Griffiths, Jack, Foley, Ioannou, & Liu, 2017; Taggart, 2022a, 2022b, 2022c). scores has arisen from the Jive verification system and was created to modularise the Jive verification functions and make them available as an open source package. scores also includes additional metrics that Jive does not contain.
scores has been used to explore user-focused approaches to evaluating probabilistic and categorical forecasts (Loveday, Taggart, & Khanarmuei, 2024).
@@ 2.4 Related Software Packages
xskillscore (Bell et al., 2021) provides many but not all of the same functions as scores and does not have direct support for pandas. The Jupyter Notebook tutorials in scores cover a wider array of metrics.
@@
Verif (Nipen, Stull, Lussana, & Seierstad, 2023) is a command line tool for generating verification plots whereas scores provides a Python API for generating numerical scores.
Pysteps (Imhoff et al., 2023; Pulkkinen et al., 2019) is a package for short-term ensemble prediction systems, and includes a significant verification submodule with many useful verification scores. PySteps does not provide a standalone verification API.

### v2

@@
scores: A Python package for verifying and evaluating models and predictions with xarray
[…]
@@ 1 Summary
scores is a Python package containing mathematical functions for the verification, evaluation and optimisation of forecasts, predictions or models. It supports labelled n-dimensional (multidimensional) data, which is used in many scientific fields and in machine learning. At present, scores primarily supports the geoscience communities; in particular, the meteorological, climatological and oceanographic communities.
scores not only includes common scores (e.g., Mean Absolute Error), it also includes novel scores not commonly found elsewhere (e.g., FIxed Risk Multicategorical (FIRM) score, Flip-Flop Index), complex scores (e.g., threshold-weighted continuous ranked probability score), and statistical tests (such as the Diebold Mariano test). It also contains isotonic regression which is becoming an increasingly important tool in forecast verification and can be used to generate stable reliability diagrams. Additionally, it provides pre-processing tools for preparing data for scores in a variety of formats including cumulative distribution functions (CDF). At the time of writing, scores includes over 50 metrics, statistical techniques and data processing tools.
@@
scores supports xarray datatypes, allowing it to work with Earth system data in a range of formats including NetCDF4, HDF5, Zarr and GRIB among others. scores uses Dask for scaling and performance. Support for pandas is being introduced.
The scores software repository can be found at https://github.com/nci/scores/.
@@ 2 Statement of Need
Labelled, n-dimensional data is widely used in many scientific fields. The Earth system science community makes heavy use of physics-based and machine learning models, both to process observations (such as identifying land use from satellite data) and to make predictions about the future (such as forecasting the weather). These models, predictions and forecasts undergo verification and evaluation to assess their correctness.
The purpose of scores is (a) to mathematically verify and validate models and predictions and (b) to foster research into new scores and metrics.
scores handles dimensionality and weighting (e.g., latitude weighting) more effectively than commonly-used data science packages. While there are existing open source Python verification packages for labelled n-dimensional data (see “Related Software Packages” further below), none of these packages offer all of the key benefits of scores.
@@ 2.1 Key Benefits of scores
To meet the needs of researchers and other users, scores provides the following key benefits.
2.1.1 Data Handling
  • Works with labelled, n-dimensional data (e.g., geospatial, vertical and temporal dimensions) for both point-based and gridded data. scores can effectively handle the dimensionality, data size and data structures commonly used for:
    • gridded Earth system data (e.g., numerical weather prediction models)
    • tabular, point, latitude/longitude or site-based data (e.g., forecasts for specific locations).
@@ 2.1.1 Data Handling
  • Supports xarray (Hoyer & Hamman, 2017) datatypes, and works with NetCDF4 (Unidata, 2024), HDF5 (The HDF Group & Koziol, 2020), Zarr (Miles et al., 2020) and GRIB (World Meteorological Organization, 2024) data formats among others.
2.1.2 Usability
@@
  • Novel scores not commonly found elsewhere (e.g., FIRM (Taggart et al., 2022), Flip-Flop Index (Griffiths et al., 2019, 2021)).
  • Commonly-used scores are also included, meeting user requests to use scores as a standalone package.
@@ 2.1.3 Compatibility
  • Support for pandas (McKinney, 2010; The pandas development team, 2024) is being introduced.
@@ 2.2 Metrics, Statistical Techniques and Data Processing Tools Included in scores
Continuous Scores for evaluating single-valued continuous forecasts. Mean Absolute Error (MAE), Mean Squared Error (MSE), Root Mean Squared Error (RMSE), Additive Bias, Multiplicative Bias, Pearson’s Correlation Coefficient, Flip-Flop Index (Griffiths et al., 2019, 2021), Quantile Loss, Murphy Score (Ehm et al., 2016).
Probability Scores for evaluating forecasts that are expressed as predictive distributions, ensembles, and probabilities of binary events. Brier Score (Brier, 1950), Continuous Ranked Probability Score (CRPS) for Cumulative Distribution Functions (CDFs) (including threshold-weighting, see Gneiting & Ranjan (2011)), CRPS for ensembles (Ferro, 2013; Gneiting & Raftery, 2007), Receiver Operating Characteristic (ROC), Isotonic Regression (reliability diagrams) (Dimitriadis et al., 2021).
Categorical Scores (including contingency table metrics) for evaluating forecasts of categories. Probability of Detection (POD), Probability of False Detection (POFD), False Alarm Ratio (FAR), Success Ratio, Accuracy, Peirce’s Skill Score (Peirce, 1884), Critical Success Index (CSI), Gilbert Skill Score (Gilbert, 1884), Heidke Skill Score, Odds Ratio, Odds Ratio Skill Score, F1 Score, Symmetric Extremal Dependence Index (Ferro & Stephenson, 2011), FIxed Risk Multicategorical (FIRM) Score (Taggart et al., 2022).
@@
Statistical Tests Tools to conduct statistical tests and generate confidence intervals. Diebold-Mariano (Diebold & Mariano, 1995) with both the Harvey et al. (1997) and Hering & Genton (2011) modifications.
@@ 2.3 Use in Academic Work
In 2015, the Australian Bureau of Meteorology began developing a new verification system called Jive, which became operational in 2022. For a description of Jive see Loveday, Griffiths, et al. (2024). The Jive verification metrics have been used to support several publications (Foley & Loveday, 2020; Griffiths et al., 2017; Taggart, 2022a, 2022b, 2022c).
scores has arisen from the Jive verification system and provides Jive verification functions as a modular, open source package. scores also includes additional metrics that Jive does not contain.
scores has been used to explore user-focused approaches to evaluating probabilistic and categorical forecasts (Loveday, Taggart, et al., 2024).
@@ 2.4 Related Software Packages
xskillscore (Bell et al., 2021) provides many but not all of the same functions as scores. The Jupyter Notebook tutorials in scores cover a wider array of metrics.
@@
Verif (Nipen et al., 2023) is a command line tool for generating verification plots whereas scores provides a Python API for generating numerical scores.
Pysteps (Imhoff et al., 2023; Pulkkinen et al., 2019) is a package for producing short-term ensemble predictions, focusing on probabilistic nowcasting of radar precipitation fields. It includes a significant verification submodule with many useful verification scores. Pysteps does not provide a standalone verification API.